You are conducting a multinomial hypothesis test (αα = 0.05) for the claim that all 5 categories are equally likely to be selected.

Category Observed
Frequency
A 17
B 11
C 24
D 25
E 25



What is the chi-square test-statistic for this data?
χ2=χ2= Incorrect

What are the degrees of freedom for this test?
d.f. = Incorrect

What is the p-value for this sample? (Report answer accurate to four decimal places.)
p-value = Incorrect
Answer #1

applying chi square goodness of fit test:

           relative observed Expected residual Chi square
category frequency Oi Ei=total*p R2i=(Oi-Ei)/√Ei R2i=(Oi-Ei)2/Ei
A 0.200 17 20.40 -0.75 0.567
B 0.200 11 20.40 -2.08 4.331
C 0.200 24 20.40 0.80 0.635
D 0.200 25 20.40 1.02 1.037
E 0.200 25 20.40 1.02 1.037
total 1.000 102 102 7.608

chi-square test-statistic X2 =7.608

degree of freedom =categories -1=5-1=4

p value =0.1070